PLEASE HELP ME SOLVE: h/2+h/5=1 and n/n+1 +3 = 4/n+1. Thank you!
Answer by checkley77(12569)

h/2+h/5=1
(5N+2H)/2*5=1
7H/10=1
7H=10
H=10/7
PROOF:
(10/7)*1/2+(10/7)*1/5=1
10/14+10/35=1
(10*35+10*14)/14*35=1
(350+140)/490=1
490/490=1
1=1
------------------------------------------
n/n+1 +3 = 4/n+1
n/(n+1)-4/(n+1)=-3
(n-4)/(n+1)=-3
n-4=-3(n+1)
n-4=-3n-3
n+3n=-3+4
4n=1
n=1/4
Proof:
(1/4)/(1/4+1)+3=4/(1/4+1)
(1/4/(5/4)+3=4/(5/4)
1/4*4/5+3=4*4/5
1/5+3=16/5
3+1/5=3+1/5